The ratio of the measures of the three angles 2:5:3

Answer by saifoo.khan(162)

Hello,
The sum of the measures of all the angles in the triangle is 180 degrees.
The ratio 2:5:3 divides the sum into 2 + 5 + 3 = 10 parts, into which the 180 degrees are divided ==>> each part = 18 degrees.
First angle: 2 parts = 2*18 = 36 degrees
Second angle: 5 parts = 5*18 = 90 degrees
Third angle: 3 parts = 3*18 = 54 degrees
Check: 36 + 90 + 54 = 180 -- yes.
Ratio 36:90:54 is 2:5:3 -- yes.
=======================
Cheers,
SAifoo.khan
einstein.saifoo@gmail.com